OpenStudy (anonymous):

Consider the equation y=-2x+5. Create a table of five ordered pairs that satisfy the equation. What is the slope of the line represented by the equation?

OpenStudy (anonymous):

is there choices

OpenStudy (anonymous):

Yes there is. A)m=2 B)m=5 C)m=-2 D)m=-5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c. (using y=mx + b formula, m=slope)

OpenStudy (anonymous):

Well, What about this? The value, in dollars, of a particular KX37B computer is given by V = -412.48x + 6843, where x is the number of years the computer has been in existence a. -412.48 signifies the amount of depreciation in one year, and 6843 signifies the initial cost. c. -412.48 signifies the hourly cost of repairs, and 6843 signifies the cost of software. b. -412.48 signifies the number of computers owned, and 6843 signifies the cost of electricity. d. 6843 signifies the amount of depreciation in one year, and -412.48 signifies the initial cost.

OpenStudy (anonymous):

i mean a or c

OpenStudy (anonymous):

i think A

OpenStudy (anonymous):

can i get a medal

OpenStudy (texaschic101):

1. C.....(y = mx + b where m is the slope) to find 5 ordered pair that will satisfy this equation, sub in any number for x and solve for y. y = -2x + 5 lets make x = 0 y = -2(0) + 5 y = 5 so when x is 0, y = 5...so one solution is : (0,5) now pick another number for x....lets let x = 1 y = -2x + 5 y = -2(1) + 5 y = -2 + 5 y = 3 so when x is 1, y is 3.....so another solution is (1,3) you can find 3 more solutions by subbing in any number and solving for y. ==================== 2. A